Johnson Products earned $4.45 per share last year and paid a dividend of $1.80 per share. If ROE was 17 percent, what is the sustainable growth rate? Do not round intermediate calculations. Enter your answer as a percent rounded to two decimal places.
Answer #1

Retention ratio = 1 - ($1.80/$4.45) = 0.59551

Sustainable growth rate = 0.17 * 0.59551 = 0.1012 or 10.12%